OpenStudy (anonymous):

A box contains four red marbles, six green marbles, and two blue marbles. What is the probability of pulling two green marbles followed by a blue marble if the marbles are pulled from the bag without replacement?

OpenStudy (countryboy15):

welcome to open study:)

OpenStudy (raffle_snaffle):

What is the total marble count?

OpenStudy (countryboy15):

what do you think the answer is and is there any choices

OpenStudy (anonymous):

i do not know. I thought it was 127/110

OpenStudy (anonymous):

Yes, there are choices. I know the answer, but I do not know how to get it.

OpenStudy (anonymous):

Actually that is the answer, i do not know how they got it

OpenStudy (countryboy15):

please put all the choices here and ill help you figure it out okay?

OpenStudy (anonymous):

1/22 5/144 127/110 1/24

OpenStudy (countryboy15):

ill be right back okay then i will try to help unless @raffle_snaffle helps you

OpenStudy (anonymous):

9/12 is not right at all

OpenStudy (raffle_snaffle):

give me a sec I am thinking. I know how to solve it.

OpenStudy (raffle_snaffle):

You sure those are your options?

OpenStudy (anonymous):

yes

OpenStudy (raffle_snaffle):

Okay I we need someone else here... @ganeshie8

OpenStudy (kittiwitti1):

Okay, I don't know if I'm doing this right, but tell me if it gets you the right answer e_o Total marbles: 4 red + 6 green + 2 blue = 12 marbles Ratio of green marbles to total marbles is 6:12, or 1:2.\[\text{green marble}\rightarrow\frac{green}{total}\rightarrow\frac{6}{12}=\frac{1}{2}\]This is for the FIRST green marble... now for the second You now have one less green marble, so it's 5 green. Also, the total has gone down to 11. The ratio is now 5:11\[\text{SECOND green marble}\rightarrow\frac{green-1}{total}\rightarrow\frac{6-1}{11}=\frac{5}{11}\]

OpenStudy (kittiwitti1):

*sorry I made an error on the last equation\[\text{Second green marble}\rightarrow\frac{green-1}{total-1}\rightarrow\frac{6-1}{12-1}=\frac{5}{11}\]Now you're down two marbles from the total, so you have 10 marbles. The amount of blue marbles hasn't changed if you haven't picked out any, so the ratio is now 2:10\[\frac{blue}{total-2}\rightarrow\frac{2}{12-2}=\frac{2}{10}=\frac{1}{5}\] I think you're supposed to do something with these ratios but I forgot e_o

OpenStudy (kittiwitti1):

The first problem in this link might help you http://www.varsitytutors.com/gre_math-help/how-to-find-the-probability-of-an-outcome

OpenStudy (kittiwitti1):

I'm aware of that, but I would rather the person come back to a solved problem than nothing. :)

OpenStudy (kittiwitti1):

Or, at least partially solved! ^_^

OpenStudy (countryboy15):

all good

OpenStudy (countryboy15):

have a good one